# Mingyu Yang, Mehdi Rezagholizadeh, Guihong Li , Vikram Appia, Emad Barsoum Advanced Micro Devices, Inc. (AMD)

### v1

@@ Abstract
With the growing demand for deploying large language models (LLMs) across diverse applications, improving their inference efficiency is crucial for sustainable and democratized access. However, retraining LLMs to meet new user-specific requirements is prohibitively expensive and environmentally unsustainable. In this work, we propose a practical and scalable alternative: composing efficient hybrid language models from existing pre-trained models. Our approach, Zebra-Llama, introduces a family of 1B, 3B, and 8B hybrid models by combining State Space Models (SSMs) and Multi-head Latent Attention (MLA) layers, using a refined initialization and post-training pipeline to efficiently transfer knowledge from pre-trained Transformers. Zebra-Llama achieves Transformer-level accuracy with near-SSM efficiency using only 7–11B training tokens (compared to trillions of tokens required for pre-training) and an 8B teacher. Moreover, Zebra-Llama dramatically reduces KV cache size—down to 3.9%, 2%, and 2.73% of the original for the 1B, 3B, and 8B variants, respectively—while preserving 100%, 100%, and $>$97% of average zero-shot performance on LM Harness tasks. Compared to models like MambaInLLaMA, X-EcoMLA, Minitron, and Llamba, Zebra-Llama consistently delivers competitive or superior accuracy while using significantly fewer tokens, smaller teachers, and vastly reduced KV cache memory. Notably, Zebra-Llama-8B surpasses Minitron-8B in few-shot accuracy by 7% while using 8×8$\times$8 × fewer training tokens, over 12×12$\times$12 × smaller KV cache, and a smaller teacher (8B vs. 15B). It also achieves 2.6×2.6$\times$2.6 ×3.8×3.8$\times$3.8 × higher throughput (tokens/s) than MambaInLlama up to a 32k context length. We will release code and model checkpoints upon acceptance.
@@ 1 Introduction
Our goal in this work is to develop a more efficient and sustainable alternative: to compose highly efficient language models directly from existing pre-trained Transformers, avoiding the cost of full pre-training while retaining performance. Our approach, called Zebra-Llama, introduces a family of hybrid models (1B, 3B, and 8B) built on two complementary components: Multi-Latent Attention (MLA) [6], a low-rank attention mechanism that compresses memory usage without sacrificing quality under moderate compression; and Mamba2, a state-space model that eliminates KV caches entirely but performs poorly when used alone [5]. Specifically, we first initialize pure MLA and pure Mamba2 models from a pre-trained Transformer via a refined weight-mapping procedure. We then use Intermediate Layer Distillation (ILD) to align their internal representations with those of the original Transformer model, ensuring strong initialization. Finally, we strategically compose hybrid architectures from the refined MLA and Mamba2 variants using a sensitivity-aware strategy called SMART (Sensitivity Measure-Aware Replacement of Transformer layers) to select where each component is most effective. This process results in highly efficient models that retain Transformer-level quality with drastically reduced memory and compute requirements. Our Zebra-Llama family of 1B, 3B, and 8B hybrid models, achieving 25×25$\times$25 ×, 50×50$\times$50 ×, and 36×36$\times$36 × KV cache compression relative to their respective base Transformer models, while maintaining 100%, 100%, and >97% of the base model’s average zero-shot performance on the LM Harness evaluation benchmark. Our models also perform competitively on few-shot tasks. Notably, Zebra-Llama-8B improves the average few-shot accuracy over Minitron-8B by 7%, despite using 8×8$\times$8 × fewer training tokens, >12$\times$ smaller KV cache, and a smaller teacher model (8B vs. 15B). Additionally, our Zebra-Llama exhibits high inference speed. Compared to the existing work MambaInLlama, it achieves 2.6–3.8$\times$ higher throughput for sequences up to 32k tokens. This work introduces a practical route for building efficient, customizable LLMs from existing models. Our key contributions are:
@@ 2 Related Work
Prior work on hybrid models can broadly be categorized into two groups based on their training strategy: pre-training-based and post-training-based approaches. Pre-training-based methods, such as Jamba [7], Hymba [8], Samba [3] and Mamba-2-Hybrid [9], interleave heterogeneous hybrid layers during full-scale model training, allowing the development of hybrid models from scratch. While these models achieve strong performance, their training cost remains high, limiting accessibility and sustainability. In contrast, post-training approaches insert efficient modules into pre-trained Transformers, often leveraging knowledge distillation to transfer capabilities without full re-training, including MambaInLLaMA [5], MOHAWK [10], and Llamba [11]. Our work focuses on the post-training setting and pushes it further by targeting extremely efficiency—both in terms of training tokens and runtime memory. We introduce a systematic method to build hybrid models with minimal training costs, making them practical for broad deployment, especially where hardware resources are critically constrained."
@@ 3.1.1 Refined Initialization
where $W^{QR}\in\mathbb{R}^{r_{q}\times n_{h}d_{r}}$ and $W^{KR}\in\mathbb{R}^{d\times d_{r}}$ The final queries and keys are then constructed as: $Q=[Q^{C};Q^{R}],\quad K=[K^{C};\text{repeat}(K^{R})]$, where $\text{repeat}(K^{R})$ denotes replication across heads. Overall, MLA reduces KV-cache size for inference from $O(n_{h}d_{h}l)$ to $O((r_{kv}+d_{r})l)$.
@@
$\begin{split}\mathcal{L}^{\text{\small{Mamba2}}}_{\text{ILD}}=\sum_{l\in\{1,2,...,L\}}\left\|h^{\text{Attn}}_{l}-h^{\text{M2}}_{l}\right\|_{2}^{2},~{}~{}\mathcal{L}^{\text{\small{MLA}}}_{\text{ILD}}=\sum_{l\in\{1,2,...,L\}}\left\|h^{\text{Attn}}_{l}-h^{\text{MLA}}_{l}\right\|_{2}^{2},\end{split}$ (11)
where $L$ denotes the total number of layers and $h_{l}$ refers to the output of the $l^{th}$ layer. Unlike MOHAWK, where the weights of MLP within each layer remain frozen during distillation, we allow training of all parameters in the Mamba2 and MLA layers. The refined initialization resulting from this ILD has proven to be crucial for enhancing the subsequent end-to-end distillation process, as evidenced in Section 4.3.3.
@@ 3.1.2 SMART: Sensitivity Measure-Aware Replacement of Transformer Layers
  • Near-Uniform Intermediate Distribution: Given the range $[L_{1}^{\text{MLA}}+1,L_{N}^{\text{MLA}}-1]$, we aim to place the remaining $N-2$ MLA layers such that the gaps between consecutive MLA layers are as uniform as possible. We constrain the gap between adjacent MLA layers to lie within the range of min: $\lfloor\frac{L_{N}^{\text{MLA}}-L_{1}^{\text{MLA}}-N+1}{N-1}\rfloor$, and max: $\lceil\frac{L_{N}^{\text{MLA}}-L_{1}^{\text{MLA}}-N+1}{N-1}\rceil$. We enumerate all valid configurations $\{C_{j}\}_{j=1}^{M}$ that meet this spacing constraint, where each candidate $C_{j}$ defines a possible set of intermediate MLA layer indices.
@@ 4.2 Performance Evaluation
Model and Setting Teacher Tokens Size KV Size ARC ARE HS MM OB PI RA WG Avg.
@@
[…]
Zebra-Llama-8B, 8MLA-24M2 8B 11B 8.38B 2.73% 57.17 79.46 77.43 55.4 44.6 79.71 44.31 71.43 63.69
@@
As shown in Table 1, compared to the base Llama models, our Zebra-Llama achieves extreme KV cache compression without noticeable performance drop. For 1B and 3B models, we achieves $3.91\%$ (25.6×25.6$\times$25.6 × compression) and $2.01\%$ (49.78×49.78$\times$49.78 × compression) KV cache size with even higher performance than the base Llama models. For the 8B models, we reach $5.47\%$ (18.3×18.3$\times$18.3 × compression) and $2.73\%$ (36.6×36.6$\times$36.6 × compression) KV cache size with only $0.87\%$ and $2.8\%$ performance drop. Note that for the 8B models, we achieve such compression ratio using only same size teachers. Further more, our Zebra-Llama achieves a better balance between KV cache compression and performance with much more efficient training (i.e., fewer training tokens or smaller teacher size) than other distillation-based methods. For example, compared with hybrid MambaInLlama, Zebra-Llama has similar performance with 12.79×12.79$\times$12.79 × and 24.88×24.$8$ smaller KV cache for 1B and 3B models. For 8B models, Zebra-Llama achieves 9.14×9.14$\times$9.14 × KV cache compression with only $1.6\%$ performance degradation by using a much smaller teacher (8B) and fewer training tokens (11B). Similarly, our method has roughly the same performance as the pure-MLA X-EcoMLA models, but with up to 3.4×3.4$\times$3.4 × less KV cache size. Moreover, compared to the pure Mamba model Llamba-8B, Zebra-Llama shows significantly better performance with both smaller teacher and fewer training tokens with minimal KV cache overhead. Moreover, we further compare Zebra-Llama with state-of-the-art hybrid models trained from scratch in Appendix C (see Table 8). While prior methods like SAMBA and Mamba-2-Hybrid rely on 1.5T–3.5T training tokens, our Zebra-Llama models achieve competitive or superior performance using only 7–11B tokens—representing a 214×–500× reduction in training data.
Few-shot Results
In Table 2, we report the results of few-shot evaluations on the Zebra-Llama-8B models under 25-shot ARC-Challenge (ARC), 10-shot HellaSwag (HS), 5-shot Winogrande (WG), 5-shot MMLU (MM), and 0-shot mc2 for Truthful-QA (TQ) tasks. Among all models, our Zebra-Llama achieves the best performance with only $5.47\%$ KV cache usage. The closest one to us is MambaInLlama-50% but it’s trained with 1.8×1.8$\times$1.8 × more tokens and has 9.14×9.14$\times$9.14 × more KV cache usage. For pure Mamba2 models such as MambaInLlama-8B-0% and Llamba-8B, they don’t use any KV cache but their performance is significantly worse than Zebra-Llama even with more training tokens and larger teachers.
Model and Setting KV % ARC HS MM WG TQ Avg.
@@
Figure 4: Inference throughput vs. output sequence length of various 8B-size models. We measure the throughput under batch size 1.
In Figure 4, we demonstrate the inference efficiency of Zebra-Llama by evaluating the throughput of different models employing distinct attention replacement strategies for Llama-3.1-8B. All experiments are conducted on a single AMD MI300X GPU. We fix the batch size and prompt length to 1 and measure the inference throughput across various generation lengths. As shown, our Zebra-Llama significantly outperforms the throughput of the MambaInLlama model (hybrid GQA-Mamba2). For long generation lengths like 32k, our method achieves throughput improvements of 2.6x and 3.79x with 8 MLA and 16 MLA layers compared to MambaInLlama with 8 GQA and 16 GQA layers. We note that MambaInLlama-8B-0% has slightly higher througput thanZebra-Llama; however, it has much worse accuracy scores accuracy, shown in Table 2.
@@ 4.3.1 Effect of Initialization Strategies
Figure 5 presents our assessment of various initialization methodologies through a comparison of three scenarios: Random weight initialization without ILD, Structured weight initialization without ILD, and the proposed Structured weight initialization with ILD. The results highlight that both structured weight initialization and ILD significantly boost SFT performance, especially when used together. For Mamba layers, ILD is crucial for aligning their outputs with the original models due to architectural differences from attention layers, providing the primary performance uplift. As for MLA layers, structured initialization can boost the accuracy significantly, which offers a strong foundation further refined by ILD.
4.3.2 Impact of SMART Layer Selection
@@
Table 3: SFT results with different layer selection strategies on Llama-3.2-1B model.
In Table 3, we demonstrate the benefits of our SMART layer selection strategy by comparing with other layer selections for our Zebra-Llama-1B with 4 MLA layers, by using the configs shown in Figure 3 as an example. First, by comparing the first two layer selections sets and our optimal solutions, we can conclude that Terminal Preservation strategy (i.e., always selecting layers from the first and last few layers) are important for preserving model accuracy. Second, we also testify the naive greed selection greedy method (the third row); the results shows that uniformly distributing the MLA layers has a major contribution to final performance as well, which verified the effectiveness of our Near-Uniform Intermediate Distribution strategy. Third, by analyzing the last three selections, together with Terminal Preservation and Near-Uniform Intermediate Distribution , Maximal Sensitivity Scores is indeed a good indicator for final accuracy score. In short, these outcomes underscore how each of the three pillars–when appropriately combined–contributes critically to the SMART layer selection strategy’s effectiveness.
@@ 4.3.3 Tradeoff between number of MLA layers and $r_{kv}$
The KV-cache size in Zebra-Llama is determined by two factors: the number of MLA layers and the KV rank $r_{kv}$ of each MLA layer. Figure 6 presents our findings from varying these two factors for our 1B model while maintaining a constant total KV-cache size. We observed that for significant KV-cache compressions, such as 12.8×12.8$\times$12.8 × and 25.6×25.6$\times$25.6 ×, optimal performance typically occurs with a moderate number of MLA layers, around six, coupled with an intermediate $r_{kv}$. Deviating from this balance will hurt the model’s performance. On one hand, increasing the number of MLA layers excessively, like to 16 layers at a 25.6×25.6$\times$25.6 × compression, forces the $r_{kv}$ per layer to become very small (e.g., $r_{kv}$ =8), which significantly degrades performance. On the other hand, reducing the number of MLA layers too much also leads to a decline in performance, since the model will then consist almost entirely of Mamba2 layers, which typically have a higher capacity gap with MHA layers than MLA layers do.

### v2

@@ Abstract
With the growing demand for deploying large language models (LLMs) across diverse applications, improving their inference efficiency is crucial for sustainable and democratized access. However, retraining LLMs to meet new user-specific requirements is prohibitively expensive and environmentally unsustainable. In this work, we propose a practical and scalable alternative: composing efficient hybrid language models from existing pre-trained models. Our approach, Zebra-Llama, introduces a family of 1B, 3B, and 8B hybrid models by combining State Space Models (SSMs) and Multi-head Latent Attention (MLA) layers, using a refined initialization and post-training pipeline to efficiently transfer knowledge from pre-trained Transformers. Zebra-Llama achieves Transformer-level accuracy with near-SSM efficiency using only 7–11B training tokens (compared to trillions of tokens required for pre-training) and an 8B teacher. Moreover, Zebra-Llama dramatically reduces KV cache size—down to 3.9%, 2%, and 2.73% of the original for the 1B, 3B, and 8B variants, respectively—while preserving 100%, 100%, and >>97% of average zero-shot performance on LM Harness tasks. Compared to models like MambaInLLaMA, X-EcoMLA, Minitron, and Llamba, Zebra-Llama consistently delivers competitive or superior accuracy while using significantly fewer tokens, smaller teachers, and vastly reduced KV cache memory. Notably, Zebra-Llama-8B surpasses Minitron-8B in few-shot accuracy by 7% while using $8\times$ fewer training tokens, over $12\times$ smaller KV cache, and a smaller teacher (8B vs. 15B). It also achieves $1.4\times$$3.3\times$ higher throughput (tokens/s) than MambaInLlama. The source code is released at https://github.com/AMD-AGI/AMD-Hybrid-Models.
@@ 1 Introduction
[…]
@@ 2 Related Work
Prior work on hybrid models can broadly be categorized into two groups based on their training strategy: pre-training-based and post-training-based approaches. Pre-training-based methods, such as Jamba [20], Hymba [8], Samba [29] and Mamba-2-Hybrid [32], interleave heterogeneous hybrid layers during full-scale model training, allowing the development of hybrid models from scratch. While these models achieve strong performance, their training cost remains high, limiting accessibility and sustainability. In contrast, post-training approaches insert efficient modules into pre-trained Transformers, often leveraging knowledge distillation to transfer capabilities without full re-training, including MambaInLLaMA [33], MOHAWK [2], and Llamba [1]. Our work focuses on the post-training setting and pushes it further by targeting extremely efficiency—both in terms of training tokens and runtime memory. We introduce a systematic method to build hybrid models with minimal training costs, making them practical for broad deployment, especially where hardware resources are critically constrained.
@@ 3.1.1 Refined Initialization
where $W^{QR}\in\mathbb{R}^{r_{q}\times n_{h}d_{r}}$ and $W^{KR}\in\mathbb{R}^{d\times d_{r}}$ . The final queries and keys are then constructed as: $Q=[Q^{C};Q^{R}],\quad K=[K^{C};\text{repeat}(K^{R})]$, where $\text{repeat}(K^{R})$ denotes replication across heads. Overall, MLA reduces KV-cache size for inference from $O(n_{h}d_{h}l)$ to $O((r_{kv}+d_{r})l)$.
@@
$\small\begin{split}\mathcal{L}^{\text{\small{Mamba2}}}_{\text{ILD}}=\sum_{l\in\{1,2,...,L\}}\left\|h^{\text{Attn}}_{l}-h^{\text{M2}}_{l}\right\|_{2}^{2},~~\mathcal{L}^{\text{\small{MLA}}}_{\text{ILD}}=\sum_{l\in\{1,2,...,L\}}\left\|h^{\text{Attn}}_{l}-h^{\text{MLA}}_{l}\right\|_{2}^{2},\end{split}$ (11)
where $L$ denotes the total number of layers and $h_{l}$ refers to the output of the $l^{th}$ layer. Unlike MOHAWK, where the weights of MLP within each layer remain frozen during distillation, we allow training of all parameters in the Mamba2 and MLA layers. The refined initialization resulting from this ILD has proven to be crucial for enhancing the subsequent end-to-end distillation process, as evidenced in Section LABEL:ablation:ILD.
@@ 3.1.2 SMART: Sensitivity Measure-Aware Replacement of Transformer Layers
  • Near-Uniform Intermediate Distribution: Given the range $L_{1}^{\text{MLA}}+1,L_{N}^{\text{MLA}}-1]$, we aim to place the remaining N2N-2 MLA layers such that the gaps between consecutive MLA layers are as uniform as possible. We constrain the gap between adjacent MLA layers to lie within the range of min: $\lfloor\frac{L_{N}^{\text{MLA}}-L_{1}^{\text{MLA}}-N+1}{N-1}\rfloor$, and max: $\lceil\frac{L_{N}^{\text{MLA}}-L_{1}^{\text{MLA}}-N+1}{N-1}\rceil$. We enumerate all valid configurations $\{C_{j}\}_{j=1}^{M}$ that meet this spacing constraint, where each candidate $C_{j}$ defines a possible set of intermediate MLA layer indices.
@@ 4.2 Performance Evaluation
Model and Setting Teacher Tokens Size KV Size ARC ARE HS $M$ OB PI RA WG Avg.
@@
Zebra-Llama-8B, 16MLA-16M2 8B 11B 8.19B 5.47% 58.62 78.37 79.27 58.17 43.4 80.03 49.28 72.61 64.97
Zebra-Llama-8B, 8MLA-24M2 8B 11B 8.38B 2.73% 58.87 79.17 78.4 54.6 43.6 79.43 46.22 72.45 64.12
@@
As shown in Table 1, compared to the base Llama models, our Zebra-Llama achieves extreme KV cache compression without noticeable performance drop. For 1B and 3B models, we achieves $3.91\%$ ($25.6\times$ compression) and $2.01\%$ ($49.78\times$ compression) KV cache size with even higher performance than the base Llama models. For the 8B models, we reach $5.47\%$ ($18.3\times$ compression) and $2.73\%$ ($36.6\times$ compression) KV cache size with only $1\%$ and $2.15\%$ performance drop. Note that for the 8B models, we achieve such compression ratio using only same size teachers. Further more, our Zebra-Llama achieves a better balance between KV cache compression and performance with much more efficient training (i.e., fewer training tokens or smaller teacher size) than other distillation-based methods. For example, compared with hybrid MambaInLlama, Zebra-Llama has similar performance with $12.79\times$ and $24.88\times$ smaller KV cache for 1B and 3B models. For 8B models, Zebra-Llama achieves $9.14\times$ KV cache compression with only $1.6\%$ performance degradation by using a much smaller teacher (8B) and fewer training tokens (11B). Similarly, our method has roughly the same performance as the pure-MLA X-EcoMLA models, but with up to $3.4\times$ less KV cache size. Moreover, compared to the pure Mamba model Llamba-8B, Zebra-Llama shows significantly better performance with both smaller teacher and fewer training tokens with minimal KV cache overhead. Moreover, we further compare Zebra-Llama with state-of-the-art hybrid models trained from scratch in Appendix C (see Table 11). While prior methods like SAMBA and Mamba-2-Hybrid rely on 1.5T–3.5T training tokens, our Zebra-Llama models achieve competitive or superior performance using only 7–11B tokens—representing a 214×–500× reduction in training data.
Model and Setting KV % ARC HS $M$ WG TQ Avg.
@@
Few-shot Results
In Table 2, we report the results of few-shot evaluations on the Zebra-Llama-8B models under 25-shot ARC-Challenge (ARC), 10-shot HellaSwag (HS), 5-shot Winogrande (WG), 5-shot MMLU (MM), and 0-shot mc2 for Truthful-QA (TQ) tasks. Among all models, our Zebra-Llama achieves the best performance with only $5.47\%$ KV cache usage. The closest one to us is MambaInLlama-50% but it’s trained with $1.8\times$ more tokens and has $9.14\times$ more KV cache usage. For pure Mamba2 models such as MambaInLlama-8B-0% and Llamba-8B, they don’t use any KV cache but their performance is significantly worse than Zebra-Llama even with more training tokens and larger teachers. Complete few-shot results for Zebra-Llama-1B and Zebra-Llama-3B can be found in Appendix E.
Results on GSM8k
In Table 3, we evaluate the math-reasoning ability of Zebra-Llama via 8-shot GSM8K task, comparing its performance against baseline models. For the Zebra-Llama-1B architecture, the 8MLA-8Mamba2 configuration achieves the highest accuracy, even outperforming the base Llama-3.2-1B-Instruct model by 6.7%. Increasing the number of MLA layers to 12 results in a performance drop from $41.09\%$ to $29.57\%$; however, this result remains superior to both MambaInLlama-1B and Llamba-1B by $4.63\%$ and $6.75\%$, respectively. For Zebra-Llama-3B, it is slightly worse than X-EcoMLA-3B given more than 2x smaller KV cache. However, our Zebra-Llama is still 9.71% and 15.01% better than MambaInLlama-3B and Llamba-3B with 14 MLA layers and 14 Mamba2 layers. Even we decrease the number of MLA layers to 8, which leads to only 2% of the original KV cache, our Zebra-Llama is still 2.1% and 7.43% better than MambaInLlama-3B and Llamba-3B. This consistent superiority demonstrates that our hybrid model structure and tailored training strategy effectively preserve, and in many cases enhance, the math-reasoning capabilities of the underlying base model while maintaining high KV cache efficiency.
Figure 4: Inference throughput vs. context length of various 8B-size models. We measure the throughput under batch size 48 and output length 1024.
Figure 5: Inference peak memory vs. context length of various 8B-size models. The out-of-memory scenarios are marked with ’OOM’.
@@
Model KV % Tokens GSM8K (8)
Llama-3.2-1B-Instruct 100% 9T 38.51
MambaInLlama-1B 50% 7B 24.94
X-EcoMLA-1B 9.37% 7B 38.06
Llamba-1B 0% 8B 22.82
Zebra-Llama-1B (8-8) 7.81% 7B 41.09
Zebra-Llama-1B (4-12) 3.91% 7B 29.57
Llama-3.2-3B-Instruct 100% 9T 70.89
MambaInLlama-3B 50% 20B 53.06
X-EcoMLA-3B 9.37% 7B 65.28
Llamba-3B 0% 10B 47.76
Zebra-Llama-3B (14-14) 4.69% 9B 62.77
Zebra-Llama-3B (6-22) 2.01% 9B 55.19
Table 3: 8-shot GSK8K results for Zebra-Llama and baselines.
[…]
@@ 4.3.1 Effect of Initialization Strategies
Table 4: SFT results with different layer selection strategies on Llama-3.2-1B model. Except for the selected MLA indices, all models are trained with same configuration.
Figure 7 presents our assessment of various initialization methodologies through a comparison of three scenarios: Random weight initialization without ILD, Structured weight initialization without ILD, and the proposed Structured weight initialization with ILD.
The results highlight that both structured weight initialization and ILD significantly boost SFT performance, especially when used together. For Mamba layers, ILD is crucial for aligning their outputs with the original models due to architectural differences from attention layers, providing the primary performance uplift. As for MLA layers, structured initialization can boost the accuracy significantly, which offers a strong foundation further refined by ILD.
4.3.2 Impact of SMART Layer Selection
In Table 4, we demonstrate the benefits of our SMART layer selection strategy by comparing with other layer selections for our Zebra-Llama-1B with 4 MLA layers, by using the configs shown in Figure 3 as an example. First, by comparing the first two layer selections sets and our optimal solutions, we can conclude that Terminal Preservation strategy (i.e., always selecting layers from the first and last few layers) are important for preserving model accuracy. Second, we also testify the naive greed selection greedy method (the third row); the results shows that uniformly distributing the MLA layers has a major contribution to final performance as well, which verified the effectiveness of our Near-Uniform Intermediate Distribution strategy. Third, by analyzing the last three selections, together with Terminal Preservation and Near-Uniform Intermediate Distribution , Maximal Sensitivity Scores is indeed a good indicator for final accuracy score. In short, these outcomes underscore how each of the three pillars–when appropriately combined–contributes critically to the SMART layer selection strategy’s effectiveness. While we believe that SMART offers a systematic and principled alternative to heuristic layer selection, we look forward to further exploration and advancement of this direction in future work.
@@ 4.3.3 Tradeoff between number of MLA layers and $r_{kv}$
The KV-cache size in Zebra-Llama is determined by two factors: the number of MLA layers and the KV rank $r_{kv}$ of each MLA layer. Figure 7 presents our findings from varying these two factors for our 1B model while maintaining a constant total KV-cache size. We observed that for significant KV-cache compressions, such as $12.8\times$ and $25.6\times$, optimal performance typically occurs with a moderate number of MLA layers, around six, coupled with an intermediate $r_{kv}$. Deviating from this balance will hurt the model’s performance. On one hand, increasing the number of MLA layers excessively, like to 16 layers at a $25.6\times$ compression, forces the $r_{kv}$ per layer to become very small (e.g., $r_{kv}$ =8), which significantly degrades performance. On the other hand, reducing the number of MLA layers too much also leads to a decline in performance, since the model will then consist almost entirely of Mamba2 layers, which typically have a higher capacity gap with MHA layers than MLA layers do.